Determine the coordinate of the center and the length of the radius.



X²+(y-2)²=10

1
Expert's answer
2022-05-27T13:52:59-0400

The general equation of a circle is (x – h)2 + (y – k)2 = r2, where (h, k) represents the location of the circle's center, and r represents the length of its radius. 

In this case (x – 0)2 + (y – 2)2 = "(\sqrt{10})^2"

So the coordinate of the center is (0; 2) and the length of the radius is "\sqrt{10}" .